### v1

@@ 2.2. Sobolev spaces on $\Gamma$
which are respectively closed subspaces of $\mathrm{L}^{p}(\Gamma)$, $\mathrm{W}^{m,p}(\Gamma)$ and $\mathbf{W}^{m,p}(\Gamma)$.
The following two results relate Sobolev spaces on $\Gamma$ with their usual flat-domain counterparts. We omit their proofs, which follow from standard arguments.
Lemma 2.3 (product rule on $\mathrm{W}^{m,p}(\Gamma)$).
Assume $\Gamma$ is of class $C^{\max\{m,2\}}$. Then for each $\theta\in C^{m}(\Gamma)$ and $u\in\mathrm{W}^{m,p}(\Gamma)$, the product $\theta u$ belongs to $\mathrm{W}^{m,p}(\Gamma)$ and
$\nabla_{M}(\theta u)=\theta\nabla_{M}u+u\nabla_{M}\theta.$
Theorem 2.4 (norm equivalence).
According to the the notation from Section 2.1, if $u\in\mathrm{W}^{m,p}(\Gamma)$ is supported on $\mathcal{W}\cap\Gamma$, then $\widetilde{u}\in\mathrm{W}^{m,p}_{0}(\mathcal{V})$ and there exist positive constants $C_{1}$ and $C_{2}$ depending only on $\Gamma$ and $m$ such that
$C_{1}\|u\|_{m,p;\Gamma}\leq\|\widetilde{u}\|_{m,p;\mathcal{V}}\leq C_{2}\|u\|_{m,p;\Gamma}.$ (2.9)
A direct consequence of 2.4 is the following density result.
Theorem 2.5 (density).
If $\Gamma$ is of class $C^{\max\{m,2\}}$, then $C^{\max\{m,2\}}(\Gamma)$ is dense in $\mathrm{W}^{m,p}(\Gamma)$.
By using appropriate cutoff functions, it follows from 2.3 and Theorems 2.4 and 2.5 that for each $u\in\mathrm{W}^{m,p}(\Gamma)$, its $m$-fold weak extrinsic gradient $\nabla_{M}^{k}u$ coincides with the formula valid for $C^{m}(\Gamma)$ functions that results from iterating (2.6).
For the rest of this section $p\in[1,\infty)$ and $m\in\mathrm{N}$ are arbitrary unless stated otherwise, and whenever we mention the space $\mathrm{W}^{m,p}(\Gamma)$ we implicitly assume $\Gamma$ is $C^{\max\{m,2\}}(\Gamma)$.
@@
Proposition 2.6 (Gagliardo–Nirenberg–Sobolev inequality [39, §4.3]).
@@
As a consequence, if $kp<d$ and $1\leq q^{\prime}\leq\frac{dp}{d-kp}$, then $\mathrm{W}^{k,p}(\Gamma)$ is continuously embedded in $\mathrm{L}^{q^{\prime}}(\Gamma)$.
Morrey’s inequality is also valid on $\Gamma$. We refer to [39, p. 389] for the precise definition of the Hölder space $C^{0,\alpha}(\Gamma)\subseteq C(\Gamma)$ ($0<\alpha<1$).
Proposition 2.7 (Morrey’s inequality [39, Theorem 6.2]).
@@
Proposition 2.8 (compact Sobolev embedding [39, Theorem 8.1]).
@@
Proposition 2.9 (compact embedding in Hölder spaces [39, Theorem 8.2]).
@@
Since $\frac{dp}{d-p}\to\infty$ as $p\to d$, we have from 2.8 and 2.9 that $\mathrm{W}^{1,p}(\Gamma)$ is compactly embedded in $\mathrm{L}^{p}(\Gamma)$ for all $1\leq p<\infty$.
The following density result (a special case of [51, Lemma 3.7]) will prove useful.
Lemma 2.11 (density).
Let $1<p<\infty$. If $\Gamma$ is of class $C^{\max\{m,2\}}$ for some integer $m\geq 1$, then for each integer $0\leq k\leq m-1$, the space of everywhere tangential $C^{m-1}$-vector fields $\mathbf{C}^{m-1}_{t}(\Gamma)$ is dense in $\mathbf{W}^{k,p}_{t}(\Gamma):=\{\mathbf{v}\in\mathbf{W}^{k,p}(\Gamma):\text{$\mathbf{v}\cdot\boldsymbol{\nu}=0$ a.e.~on $\Gamma$}\}$.
Remark 2.12 (tangential components).
Provided $1<p<\infty$ and $\Gamma$ is of class $C^{m+1}$ for some integer $m\geq 0$, the product rule 2.3 implies: for each $\mathbf{v}\in\mathbf{W}^{m,p}_{t}(\Gamma):=\mathrm{W}^{m,p}_{t}(\Gamma,\mathrm{R}^{d+1})$ supported on $\mathcal{W}\cap\Gamma$, there exists a unique $\widehat{\mathbf{v}}=(\widehat{v}_{i})_{i=1}^{d}\in\mathbf{W}^{m,p}(\mathcal{V}):=\mathrm{W}^{m,p}(\mathcal{V},\mathrm{R}^{d})$ supported on $\mathcal{X}:=\widetilde{\boldsymbol{\chi}}^{-1}(\mathcal{W}\cap\Gamma)\subseteq\mathcal{V}$ such that
$\widetilde{\mathbf{v}}=(\nabla\widetilde{\boldsymbol{\chi}})\widehat{\mathbf{v}}\quad\text{on $\mathcal{V}$},$ (2.10)
where $\widetilde{\mathbf{v}}=\mathbf{v}\circ\widetilde{\boldsymbol{\chi}}\in\mathrm{W}^{m,p}(\mathcal{V},\mathrm{R}^{d+1})$; equivalently, $\widehat{\mathbf{v}}=\mathbf{g}^{-1}(\nabla^{T}\widetilde{\boldsymbol{\chi}})\widetilde{\mathbf{v}}$ on $\mathcal{V}$. Moreover, because of 2.4 (norm equivalence) and the Leibniz rule for Sobolev spaces in flat domains,
$\|\widehat{\mathbf{v}}\|_{m,p;\mathcal{X}}\cong\|\widetilde{\mathbf{v}}\|_{\mathrm{W}^{m,p}(\mathcal{X},\mathrm{R}^{d+1})}\cong\|\mathbf{v}\|_{m,p;\Gamma}.$ (2.11)
where the implicit equivalence constants depend only on $\Gamma$.
3. Laplace–Beltrami and Bochner–Laplace problems
@@
As was stated in the introduction, the nonlinear nature of the tangent Navier–Stokes equations necessitates studying the tangent Stokes equations in the $\mathrm{L}^{p}$-based setting. It turns out that to develop the $\mathrm{L}^{p}$-based theory for the latter, it suffices to have the $\mathrm{L}^{p}$-based theory for the Laplace–Beltrami and the Bochner-Laplace problems at hand. We illustrate this idea by the following formal calculation.
@@
Consider the following weak formulation of the tangent Stokes problem with $p\in(1,\infty)$ and $p^{*}:=\frac{p}{p-1}$: given $(\mathbf{f},g)\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}\times\mathrm{L}^{p^{*}}_{\#}(\Gamma)$, find $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ such that
@@
Using the symmetry of the covariant Hessian $\nabla_{\Gamma}\nabla_{\Gamma}\psi$ and $\operatorname{div}_{\Gamma}\mathbf{u}=g$ one can rewrite the above as
@@
If we had the $\mathrm{L}^{p}$-based regularity and well-posedness in $\mathrm{W}^{1,p}(\Gamma)$ for the Laplace–Beltrami problem for $\Gamma$ of class $C^{3}$, then we could infer that
@@
where we have used the well-posedness of problem (3.1) to bound $\|\mathbf{u}\|_{0,p;\Gamma}$ by data. With this at hand, we can now go back to the momentum equation to obtain a Bochner-Laplace equation for the velocity:
@@
If we had the $\mathbf{W}^{2,p}(\Gamma)$-regularity for the Bochner-Laplace problem, we could conclude

### v2

@@ 2.1. Differential geometry and Sobolev spaces on $\Gamma$
which are respectively closed subspaces of $\mathrm{W}^{m,p}(\Gamma)$ and $\mathbf{W}^{m,p}(\Gamma)$. Notice that, if $\Gamma$ is of class $C^{m,1}$ for some nonnegative integer $m$, $\mathrm{W}^{m+1,p}(\Gamma)\subseteq\mathrm{W}^{m,p}(\Gamma)$ and $\nabla_{M}u\in\mathbf{W}^{m,p}_{t}(\Gamma)$ for each $u\in\mathrm{W}^{m+1,p}(\Gamma)$. The following density result, which is a slight improvement over [45, Lemma 3.7]), will be instrumental in this work.
Lemma 2.7 (density).
Let $p\in[1,\infty)$, $m\in\mathrm{N}$ and an integer $0\leq k\leq m-1$. If $\Gamma$ is of class $C^{m}$, then the space of everywhere tangential $C^{m-1}$-vector fields $\mathbf{C}^{m-1}_{t}(\Gamma)$ is dense in $\mathbf{W}^{k,p}_{t}(\Gamma)$. Similarly, if $m\geq 2$ and $\Gamma$ is of class $C^{m-1,1}$, then the space of a.e. tangential $C^{m-2,1}$-vector fields $\mathbf{C}^{m-2,1}_{t}(\Gamma)$ is dense in $\mathbf{W}^{k,p}_{t}(\Gamma)$.
If $\Gamma$ is of class $C^{2}$, from [36, eq. (2.17)] and 2.7 we know that
$\mathbf{B}\mathbf{v}=-(\nabla_{M}^{T}\mathbf{v})\boldsymbol{\nu},\qquad\forall\mathbf{v}\in\mathbf{C}^{1}_{t}(\Gamma).$ (2.6)
$\nabla_{M}\mathbf{v}=\nabla_{\Gamma}\mathbf{v}-\boldsymbol{\nu}\mathbf{v}^{T}\mathbf{B},\qquad\forall\mathbf{v}\in\mathbf{C}^{1}_{t}(\Gamma).$ (2.7)
This shows that for tangential vector fields, $\nabla_{M}$ and $\nabla_{\Gamma}$ only differ in lower order terms.
For the rest of this section $p\in[1,\infty)$ and $m\in\mathrm{N}$ are arbitrary unless stated otherwise, and whenever we mention the space $\mathrm{W}^{m,p}(\Gamma)$ we implicitly assume $\Gamma$ is of class $C^{m-1,1}$.
@@
Proposition 2.8 (Gagliardo–Nirenberg–Sobolev inequality [35, §4.3]).
@@
As a consequence, if $kp<d$ and $1\leq r\leq\frac{dp}{d-kp}$, then $\mathrm{W}^{k,p}(\Gamma)$ is continuously embedded in $\mathrm{L}^{r}(\Gamma)$.
Morrey’s inequality is also valid on $\Gamma$. We refer to [35, p. 389] for the precise definition of the Hölder space $C^{0,\alpha}(\Gamma)\subseteq C(\Gamma)$ ($0<\alpha<1$).
Proposition 2.9 (Morrey’s inequality [35, Theorem 6.2]).
@@
Proposition 2.10 (compact Sobolev embedding [35, Theorem 8.1]).
@@
Proposition 2.11 (compact embedding in Hölder spaces [35, Theorem 8.2]).
@@
Since $\frac{dp}{d-p}\to\infty$ as $p\to d$, we have from 2.10 and 2.11 that $\mathrm{W}^{1,p}(\Gamma)$ is compactly embedded in $\mathrm{L}^{p}(\Gamma)$ for all $1\leq p<\infty$.
Recall the following identity valid for $\Gamma$ of class $C^{3}$ and for each $\mathbf{v}\in\mathbf{C}^{2}_{t}(\Gamma)$ [36, Lemma 2.1]:
$\mathbf{P}\operatorname{div}_{\Gamma}\nabla^{T}_{\Gamma}\mathbf{v}=\nabla_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v}+\left(\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2}\right)\mathbf{v}.$
We will show that, loosely speaking, said identity is still valid in a weak sense for $\Gamma$ of class $C^{2}$ if we replace $\nabla_{\Gamma}^{T}\mathbf{v}$ by $\nabla_{\Gamma}\mathbf{v}$ on the left-hand side (hence obtaining the Bochner Laplacian $\Delta_{B}\mathbf{v}=\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{v}$) as long as we test against gradients of scalar functions. This elementary observation, which relies on $\Gamma$ not having a boundary, will play a crucial role throughout this paper. We refer to Appendix B for its proof.
Lemma 2.13 (weak commutator identity).
Assume $\Gamma$ is of class $C^{2}$. Then for each $\phi\in C^{2}(\Gamma)$ and $\mathbf{v}\in\mathbf{C}^{1}_{t}(\Gamma)$ there holds that
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{v}:\nabla_{\Gamma}\nabla_{\Gamma}\phi=-\int_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v}\,\Delta_{\Gamma}\phi+\int_{\Gamma}\left(\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2}\right)\mathbf{v}\cdot\nabla_{\Gamma}\phi.$ (2.8)
2.2. The Laplace-Beltrami problem
In this section, we summarize some of our main results concerning the Laplace–Beltrami problem posed on $\Gamma$ from the companion paper [5]. We begin with the strong formulation:
$\Delta_{\Gamma}u=f,$ (LB)
where $\Delta_{\Gamma}$ is the Laplace–Beltrami operator $\Delta_{\Gamma}=\operatorname{div}_{\Gamma}\nabla_{\Gamma}$. Before we begin, the following result will be of use in the forthcoming analysis.
Proposition 2.14 (surjectivity of the divergence operator).
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{0,1}$. Then, the linear and continuous operator $\operatorname{div}_{\Gamma}:\mathbf{L}^{p}_{t}(\Gamma)\rightarrow(\mathrm{W}^{1,p^{*}}_{\#}(\Gamma))^{\prime}$ defined by
$\langle\operatorname{div}_{\Gamma}\mathbf{u},v\rangle=-\int_{\Gamma}\mathbf{u}\cdot\nabla_{\Gamma}v,\qquad\forall\mathbf{u}\in\mathbf{L}^{p}_{t}(\Gamma),\ \forall v\in\mathrm{W}^{1,p^{*}}_{\#}(\Gamma),$
is surjective.
We now state the main result from [5] concerning well-posedness and higher regularity of solutions to the Laplace–Beltrami problem.
Theorem 2.15 (Laplace–Beltrami).
Let $p\in(1,\infty)$ and assume that $\Gamma$ is of class $C^{1}$ or of class $C^{0,1}$ if $p\in(2-\varepsilon,2+\varepsilon)$ for a certain $\varepsilon>0$ uniquely determined by $\Gamma$. Then, for every $f\in(\mathrm{W}^{1,p^{*}}_{\#}(\Gamma))^{\prime}$, there exists a unique $u\in\mathrm{W}^{1,p}_{\#}(\Gamma)$ such that
Moreover, there exists a constant $C>0$, depending only on $\Gamma$ and $p$, such that
Furthermore, if $\Gamma$ is of class $C^{m+1,1}$ for some nonnegative integer $m$ and $f\in\mathrm{W}^{m,p}_{\#}(\Gamma)$, then the solution $u\in\mathrm{W}^{1,p}_{\#}(\Gamma)$ of (2.9) is in fact in $\mathrm{W}^{m+2,p}_{\#}(\Gamma)$ and satisfies the PDE in strong form: $\Delta_{\Gamma}u=f$ a.e. on $\Gamma$. Moreover, there is $C=C(\Gamma,p)>0$ such that
$\|u\|_{m+2,p;\Gamma}\leq C\|f\|_{m,p;\Gamma}.$
The next corollary will be of use when analyzing the Bochner–Laplace problem (VL).
Corollary 2.16.
Let $p\in(1,\infty)$ and $q\in(1,d)$, and define $s:=\min\{p,\frac{dq}{d-q}\}$. Assume $\Gamma$ is of class $C^{1}$ or of class $C^{0,1}$ if $p\in(2-\varepsilon,2+\varepsilon)$ for a certain $\varepsilon>0$ uniquely determined by $\Gamma$. Then for each $\mathbf{f}\in\mathbf{L}^{p}_{t}(\Gamma)$ and $f\in\mathrm{L}^{q}_{\#}(\Gamma)$, there is a unique $u\in\mathrm{W}^{1,s}_{\#}(\Gamma)$ such that
$\int_{\Gamma}\nabla_{\Gamma}u\cdot\nabla_{\Gamma}v=-\int_{\Gamma}\mathbf{f}\cdot\nabla_{\Gamma}v+\int_{\Gamma}f\,v,\qquad\forall v\in\mathrm{W}^{1,s^{*}}_{\#}(\Gamma).$
Moreover, there exists a positive constant $C$, depending only on $\Gamma$, $p$ and $q$ such that
$\|u\|_{1,s;\Gamma}\leq C\left(\|\mathbf{f}\|_{0,p;\Gamma}+\|f\|_{0,q;\Gamma}\right)$
We close this section with the discussion of “ultra-weak” solutions to (LB), which will be utilized in the analyses of (VL) and (S).
Lemma 2.17 (ultra-weak solutions of the Laplace–Beltrami operator on $\Gamma$).
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{1,1}$. Then, for each $f\in(\mathrm{W}^{2,p^{*}}_{\#}(\Gamma))^{\prime}$ there exists a unique $u\in\mathrm{L}^{p}_{\#}(\Gamma)$ such that
Moreover, there exists a constant $C>0$ depending only on $\Gamma$ and $p$ such that
$\|u\|_{0,p;\Gamma}\leq C\|f\|_{(\mathrm{W}^{2,p^{*}}_{\#}(\Gamma))^{\prime}}.$
3. The Bochner–Laplace problem
@@
As was stated in the introduction, the nonlinear nature of the tangent Navier–Stokes equations necessitates studying the tangent Stokes equations in the $\mathrm{L}^{p}$-based setting. It turns out that to develop the $\mathrm{L}^{p}$-based theory for the latter, it suffices to have the $\mathrm{L}^{p}$-based theory for the Laplace–Beltrami and the Bochner–Laplace problems at hand. We illustrate this idea by the formal calculation below.
@@
Consider the weak formulation of the tangent Stokes problem with $p\in(1,\infty)$: given $(\mathbf{f},g)\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}\times\mathrm{L}^{p}_{\#}(\Gamma)$, find $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ such that
@@
Using the symmetry of the covariant Hessian (cf. (2.2)) $\nabla_{\Gamma}\nabla_{\Gamma}\psi$ and $\operatorname{div}_{\Gamma}\mathbf{u}=g$, in conjunction with (2.8), one can rewrite the above as
@@
If we had the $\mathrm{L}^{p}$-based regularity and well-posedness in $\mathrm{W}^{1,p}(\Gamma)$ for the Laplace–Beltrami problem for $\Gamma$, then we could infer that
@@
where we have used the well-posedness of problem (3.1) to bound $\|\mathbf{u}\|_{0,p;\Gamma}$ by data. With this at hand, we can now go back to the momentum equation to obtain a Bochner–Laplace equation for the velocity:
@@
If we had the $\mathbf{W}^{2,p}(\Gamma)$-regularity for the Bochner–Laplace problem, we could conclude